OpenStudy (anonymous):

The length of a rectangle is 5 inches more than three times the width. The perimeter is 90 inches. Find the length and width

OpenStudy (anonymous):

L=length W=width The relation --> L=3W+5 2L+2W=perimeter of the rectangle >now substitute 3W+5 for L, so that you are dealing with like terms 2(3W+5)+2W=permiteter >we already know that the perimeter is 90, so make that equal to the above expression Step by step calculation: 2(3W+5)+2W=90 6W+10+2W=90 8W+10=90 8W=80 *w=10 now you can sub W=10 into the original relation for L L=3W+5 L=3(10)+5 L=30+5 *L=35